# Accelerating ViT Inference on FPGA through Static and Dynamic Pruning

### 21 Mar 2024 v1

@@ Abstract
Vision Transformers (ViTs) have achieved state-of-the-art accuracy on various computer vision tasks. However, their high computational complexity prevents them from being applied to many real-world applications. Weight and token pruning are two well-known methods for reducing complexity–weight pruning reduces the model size and associated computational demands, while token pruning further dynamically reduces the computation based on the input. Combining these two techniques should significantly reduce computation complexity and model size; however, naively integrating them results in irregular computation patterns, leading to significant accuracy drops and difficulties in hardware acceleration.
Addressing the above challenges, we propose a comprehensive algorithm-hardware codesign for accelerating ViT on FPGA through simultaneous pruning–combining static weight pruning and dynamic token pruning. For algorithm design, we systematically combine a hardware-aware structured block-pruning method for pruning model parameters and a dynamic token pruning method for removing unimportant token vectors. Moreover, we design a novel training algorithm to recover the model’s accuracy. For hardware design, we develop a novel hardware accelerator for executing the pruned model. The proposed hardware design employs multi-level parallelism with load balancing strategy to efficiently deal with the irregular computation pattern led by the two pruning approaches. Moreover, we develop an efficient hardware mechanism for efficiently executing the on-the-fly token pruning. We apply our codesign approach to the widely used DeiT-Small model. We implement the proposed accelerator on a state-of-the-art FPGA board. The evaluation results show that the proposed algorithm can reduce computation complexity by up to 3.4×3.4$\times$3.4 × with $\approx 3\%$ accuracy drop and a model compression ratio of up to 1.6×1.6$\times$1.6 ×. Compared with state-of-the-art implementation on CPU, GPU, and FPGA, our codesign on FPGA achieves average latency reduction of 12.8×12.8$\times$12.8 ×, 3.2×3.2$\times$3.2 ×, 0.72.1×0.7-2.1$\times$0.7 - 2.1 × respectively.
@@ I Introduction
Vision Transformers (ViTs) [1] have demonstrated superior performance in comparison to Convolutional Neural Networks (CNNs) in various vision tasks [2, 3, 4, 5, 6, 7, 8]. The global self-attention in ViTs leads to a reduced local, image-specific inductive bias [1]; this results in ViTs requiring larger datasets and larger model sizes [9] to perform better than CNN. The Multi-head Self-Attention (MSA) of ViTs allows them to generalize better than CNNs on larger datasets [10]. However, their computational cost is usually significantly higher than CNNs due to the MSA mechanism, which scales quadratically with the number of input tokens [11, 12]. Their intensive computational requirements emphasize the need for efficient acceleration.
@@
In addressing the computational challenge, pruning has been proven to be effective in reducing network dimensions in CNNs [13, 14, 15, 16]. However, explorations in self-attention-based pruning methods still need to be discovered [17, 18, 19]. Many existing works on efficient ViTs explored block weight pruning and token pruning as two distinct strategies. Weight pruning, introduced in [20, 21, 22, 23, 17, 17, 24, 18, 25, 25, 26] reduces the model size by pruning input parameters statically and selectively, thus feeding the neural network with sparse inputs to reduce computation. Token pruning removes tokens to reduce the computational complexity. The static approaches in [27, 28, 29, 30] drop tokens with a fixed ratio, often ignoring the redundancies between tokens; dynamic token pruning studies in [31, 32, 33] do not fully explore the token redundancies from different attention heads and simply discard non-informative tokens. [34, 35, 36, 37] dynamically reduces the number of tokens in ViT encoders during inference based on the inherent image characteristics. Moreover, only a few of these studies support efficient hardware implementations by the respective pruning algorithm. Both weight pruning and token pruning methods reduce the computational complexity independently, but the interaction between the two remains unexplored. A combined approach could bring further computational benefits. However, such integration poses two main challenges: (1) accuracy drop (algorithm level) and (2) increased computational pattern irregularities (hardware level).
@@
Many ViT acceleration works primarily focus on the CPU and GPU platforms [28, 36, 32, 29, 30]. However, the integration of block weight pruning and token pruning in ViTs effectively reduces the model size, thus making it possible to accommodate the compressed model onto FPGA. Comprehensively, we use FPGA to accelerate our pruned ViT models for these reasons: (1) FPGAs, with customized data path and on-chip memory organization, stand out as better choices than CPU/GPU to maximize the computation efficiency. (2) CPUs and GPUs cannot effectively handle the token shuffling process of our dynamic token pruning. We design a specific FPGA kernel to handle the token shuffling in the middle of model inferences. (3) CPUs and GPUs need complicated process to address work-load imbalance, whereas on FPGA, we can design customized hardware module for balancing work-load.
@@
In this paper, we propose an algorithm-hardware codesign for accelerating ViT inference. Different from existing ViT acceleration works, we utilize the combined power of static weight pruning and dynamic token pruning. We propose a simultaneous pruning algorithm to recover the model accuracy caused by two pruning approaches. Combining the two pruning approaches leads to more severe computational irregularity than applying a single pruning approach. Therefore, we develop a customized data path and memory organization on FPGA to execute the pruned model efficiently. While existing ViT accelerators on FPGA [35, 37] can handle the irregular patterns after pruning, they target either weight pruning or token pruning, but not both. Therefore, none of the existing FPGA ViT accelerators can support our integrated simultaneous pruning approach. We summarize our main contributions below:
@@
  • We propose an algorithm-hardware codesign for efficient ViT inference based on FPGA that combines parameter (static) and token (dynamic) pruning to reduce both the ViT model size and computational complexity.
@@
  • For the hardware design, we develop a novel hardware accelerator with multi-level parallelism with load balancing strategy to efficiently deal with (1) load imbalance caused by the block pruning and (2) a challenging number of tokens caused by token pruning. We also develop an efficient hardware mechanism for executing the on-the-fly token pruning algorithm.
  • We evaluate our codesign on DeiT models and deploy the proposed accelerator on a state-of-the-art FPGA board - Xilinx Alveo U250. The evaluation results show that the proposed algorithm can reduce computation complexity by up to 3.4×3.4$\times$3.4 × with $\approx 3\%$ accuracy drop and a model compression ratio of up to 1.6×1.6$\times$1.6 ×. Compare with state-of-the-art implementation on CPU, GPU, and FPGA, our codesign on FPGA achieves average latency reduction of 12.8×12.8$\times$12.8 ×, 3.2×3.2$\times$3.2 ×, 0.72.1×0.7-2.1$\times$0.7 - 2.1 × respectively.
@@ II-A Vision Transformer
ViT [1] has a stack of transformer encoders. Each encoder has a multi-head self-attention (MSA) and a multi-layer perceptron (MLP). The input image $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$ is first partitioned into $N$ patches $\mathbf{x}_{p}\in\mathbb{R}^{N\times P^{2}C}$. Each patch is flattened into a vector of length $P^{2}C$. Next, a learnable linear mapping maps each patch to a token vector of length $D$. A special parameterized token $\mathbf{x}_{\text{CLS}}$ is appended as a token vector. Then, a positional embedding $\mathbf{E}_{\text{POS}}\in\mathbb{R}^{(N+1)\times D}$ is added to input token matrix to produce $\mathbf{Z_{0}}\in\mathbb{R}^{(N+1)\times D}$ which is the input to the transformer encoder stack. For simplicity, we denote the number of input tokens to the encoder stack as $N$ instead of $N+1$ for the rest of the paper.
@@ II-B Related Work
Weight pruning: Structured and hardware-friendly model parameter pruning, used in traditional CNNs [20, 21, 22], becomes popular for ViT. [23] proposes to prune parameters across all the encoders by building upon [17] that prunes network parameters without structure. [17] introduces the notion of movement pruning that prunes parameters by associating a learnable importance score to each parameter and generating a pruning mask based on the learned scores, penalizing the inclusion of parameters. Magnitude-based approaches, on the other hand, discard parameters with large magnitudes [24]. [18] partitions a parameter matrix into blocks and prunes the rows and columns in each block by using the $l_{2}$ norms. [25] prunes entire attention heads within the MSA and neurons in each feed-forward linear layer (width pruning) facilitated via saliency scores for each head and neuron, respectively. [25] also removes entire encoders after a certain depth (depth pruning). [26] proposes a collaborative approach to optimizing ViT pruning that prunes heads and neurons; the latter neurons are pruned such that they reduce the length of each token. This method considers the collective pruning impact through an expensive approximation of the Hessian of the loss.
Token pruning: Token pruning approaches attempt to identify redundant tokens and drop them to reduce the computational footprint associated with the number of tokens. [28] proposes a static approach to token dropping that ranks the importance of tokens by the attention score of the class token with respect to each token aggregated across heads. In theory, such static approaches do not need additional training since the token-dropping module is not parameterized. In contrast to static token pruning, dynamic token pruning as employed in [36, 35, 34] adds additional model parameters that facilitate the selection of relevant/attentive tokens. [36] and [35] utilize a token selector network inserted at various depths along the original transformer network; such token selector networks are neural networks that output a decision (binary) mask to inform the retention or removal of a token. [34], on the other hand, associate a learnable score to each token and prune tokens with scores lower than a threshold.
@@ III-A Problem Definition
Our objective is to accelerate ViT on FPGA through an algorithm-hardware codesign that 1) utilizes a novel combination of model weight pruning and token pruning to reduce computation complexity (algorithm design) and 2) an efficient accelerator that explicitly accounts for the distinct and irregular access patters of the two pruning techniques for executing the combined pruned model (hardware design).
For algorithm design, the input is a ViT model denoted as $\mathcal{M}(\cdot,h_{\text{structure}},\Theta)$ where $h_{\text{structure}}$ are the model hyper-parameters, including number of encoders, number of heads, dimensions of linear layers, etc. $\Theta$ is the trainable parameters containing the weights and biases of the MSA and MLP. Our algorithm design prunes the input model $\mathcal{M}$ through (1) offline weight pruning that reduces the redundant parameters in $\Theta$, and (2) runtime token pruning that trims the number of tokens (in $h_{\text{structure}}$) in the intermediate layers according to the importance of the token. After the pruning, we obtain the pruned model denoted as $\mathcal{M}^{{}^{\prime}}(\cdot,h_{\text{structure}}^{{}^{\prime}},\Theta^{{}^{\prime}})$ where $\Theta^{{}^{\prime}}$ denotes the model parameters after weight pruning and $h_{\text{structure}}^{{}^{\prime}}$ denotes the hyperparameters after token pruning. The objective of algorithm design is to reduce the number of parameters, reduce the computation complexity, and maintain accuracy.
For hardware design, the hardware accelerator executes the pruned model $\mathcal{M}^{{}^{\prime}}(\cdot,h_{\text{structure}}^{{}^{\prime}},\Theta^{{}^{\prime}})$ where $\Theta^{{}^{\prime}}$. For executing the model, the input is an image sample $\bm{x}$ and the accelerator executes $\mathcal{M}^{{}^{\prime}}(\bm{x},h_{\text{structure}}^{{}^{\prime}},\Theta^{{}^{\prime}})$ to obtain the result. The latency is the duration from the time when the accelerator receives the input $\bm{x}$ to the time when the accelerator obtains the result.
@@ III-B Computational Complexity
The computational complexity for each operation within the MSA and MLP wihout pruning are summarized in table I. $B$ denotes the batch size.
@@ III-C Overview of Algorithm-hardware Codesign
Hardware accelerator design: At runtime, when the host process receives an input image, it sends the input image to the accelerator for inference. We employ a multi-level parallelism strategy for the proposed hardware architecture to efficiently deal with the irregular computation patterns and design a Token Dropping Hardware Module for efficient on-the-fly token dropping. See Section (Section V-C) for details.
@@ IV Pruning Algorithm
Existing works only utilize either weight pruning or token pruning. In contrast, we systematically combine the two pruning approaches with a novel training algorithm for recovering accuracy. To this end, we first introduce static weight pruning (Section IV-A) and dynamic token pruning (Section IV-B) separately and then introduce our Simultaneous Pruning algorithm to prune the input model.
@@ IV-A1 Pruning of MSA
We use $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, $\mathbf{W}_{v}$ $\in\mathbb{R}^{D\times HD^{\prime}}$ to denote the concatenation of weight matrices of all the head. For example, $\mathbf{W}_{p}\in\mathbb{R}^{D\times HD^{\prime}}\text{ }\mbox{where}\text{ }p=\{q,k,v\}$ The projection operation (Equation 5) projects the concatenated SA outputs of embedding dimension $HD^{\prime}$ back to dimension $D$ via $\mathbf{W}_{\text{proj}}\in\mathbb{R}^{HD^{\prime}\times D}$ To prune a weight matrix $\mathbf{W}\in\mathbb{R}^{M_{1}\times M_{2}}$, we define a parameterized score matrix $\mathbf{S}\in\mathbb{R}^{m\times n}$ such that, $(m,n)=(\left\lceil\frac{M_{1}}{b}\right\rceil,\left\lceil\frac{M_{2}}{b}\right\rceil)$ where $(b,b)$ is the block size. $\mathbf{S}_{ij}$ denotes the importance score of a parameter block of size $(b,b)$ in the weight matrix $\mathbf{W}$ denoted by the slice $\mathbf{W}(ib:\alpha,jb:\beta)$ where $(\alpha,\beta)=(\min(ib+b,M_{1}),\min(jb+b,M_{2}))$. $\mathbf{S}$ is used to construct a mask $\mathbf{M}\in\mathbb{R}^{M_{1}\times M_{2}}$ via the top-$k$ selection:
@@
wher $\mathbf{M}_{ij}^{\text{block}}(.)$ is a block of size $(b,b)$ in $\mathbf{M}$ corresponding to $s_{ij}$. The masked weight is generated as, $\mathbf{W(\mathbf{M})}=\mathbf{W}\odot\mathbf{M}$ where $\odot$ is the element-wise Hadamard product. The generated masked weight $\mathbf{W(\mathbf{M})}$ is used for the forward pass during training. Note that top-$k$ is the target weight blocks of interest. To compute the gradient of $\mathbf{S}$ during backward pass, a straight-through estimator (STE) [39, 40, 41] is used that neglects the gradients of $\mathbf{M}$ with respect to $\mathbf{S}$. Additionally, the pruning of $\mathbf{W}_{p}$ in row dimension and the pruning of $\mathbf{W}_{\text{proj}}$ in column dimension follows the same pattern (denoted as alternate pattern), as shown in Figure 2. For example, a head removed from $\mathbf{W}_{p}$ makes the corresponding head in $\mathbf{W}_{\text{proj}}$ redundant, and vice-versa.

### 12 Apr 2024 v2

@@ Abstract
Vision Transformers (ViTs) have achieved state-of-the-art accuracy on various computer vision tasks. However, their high computational complexity prevents them from being applied to many real-world applications. Weight and token pruning are two well-known methods for reducing computational complexity. Weight pruning reduces the model size and associated computational demands, while token pruning further reduces the computation based on the input. Combining these two techniques should significantly reduce computation complexity and model size; however, naively integrating them results in irregular computation patterns, leading to significant accuracy drops and difficulties in hardware acceleration.
To address the above challenges, we propose a comprehensive algorithm-hardware codesign for accelerating ViT on FPGA through simultaneous pruning – combining static weight pruning and dynamic token pruning. For algorithm design, we systematically combine a hardware-aware structured block-pruning method for pruning model parameters and a dynamic token pruning method for removing unimportant token vectors. Moreover, we design a novel training algorithm to recover the model’s accuracy. For hardware design, we develop a novel hardware accelerator for executing the pruned model. The proposed hardware design employs multi-level parallelism with a load-balancing strategy to efficiently deal with the irregular computation pattern led by the two pruning approaches. Moreover, we develop an efficient hardware mechanism for executing the on-the-fly token pruning. We apply our codesign approach to the widely used DeiT-Small model. We implement the proposed accelerator on a state-of-the-art FPGA board. The evaluation results show that the proposed algorithm can reduce computation complexity by up to 3.4×3.4$\times$3.4 × with $\approx 3\%$ accuracy drop and a model compression ratio of up to 1.6×1.6$\times$1.6 ×. Compared with state-of-the-art implementation on CPU, GPU, and FPGA, our codesign on FPGA achieves an average latency reduction of 12.8×12.8$\times$12.8 ×, 3.2×3.2$\times$3.2 ×, and 0.72.1×0.7-2.1$\times$0.7 - 2.1 ×, respectively.
@@ I Introduction
Vision Transformers (ViTs) [1] have demonstrated superior performance in comparison to Convolutional Neural Networks (CNNs) in various vision tasks [2, 3, 4, 5, 6, 7, 8]. The global self-attention in ViTs leads to a reduced local and image-specific inductive bias [1]; this results in ViTs requiring larger datasets and larger model sizes [9] to perform better than CNN. The Multi-head Self-Attention (MSA) of ViTs allows them to generalize better than CNNs on larger datasets [10]. However, their computational cost is usually significantly higher than CNNs due to the MSA mechanism, which scales quadratically with the number of input tokens [11, 12]. Their intensive computational requirements emphasize the need for efficient hardware acceleration.
@@
In addressing the computational challenge, pruning has been proven to be effective in reducing the computational cost of CNNs [13, 14, 15, 16]. However, explorations in self-attention-based pruning methods still need to be discovered [17, 18, 19]. Many existing works on efficient ViTs explored block weight pruning and token pruning as two distinct strategies. Weight pruning, introduced in [20, 21, 22, 23, 17, 24, 18, 25, 25, 26], reduces the model size by pruning input parameters statically and selectively, thus feeding the neural network with sparse inputs to reduce computation. Token pruning removes tokens to reduce the computational complexity. The static approaches in [27, 28, 29, 30] drop tokens with a fixed ratio, often ignoring the redundancies between tokens; dynamic token pruning studies in [31, 32, 33] do not fully explore the token redundancies from different attention heads and simply discard non-informative tokens. [34, 35, 36, 37] dynamically reduce the number of tokens in ViT encoders during inference based on the inherent image characteristics. Moreover, only a few of these studies support efficient hardware implementations by the respective pruning algorithm. Both weight pruning and token pruning methods reduce the computational complexity independently, but the interaction between the two remains unexplored. A combined approach could bring further computational benefits. However, such integration poses two main challenges: (1) accuracy drop (algorithm level) and (2) increased computational pattern irregularities (hardware level).
@@
Many ViT acceleration works primarily focus on the CPU and GPU platforms [28, 36, 32, 29, 30]. However, the integration of block weight pruning and token pruning in ViTs effectively reduces the model size, thus making it possible to accommodate the compressed model onto FPGA. Comprehensively, we use FPGA to accelerate our pruned ViT models for these reasons: (1) FPGAs, with customized data path and on-chip memory organization, stand out as better choices than CPU/GPU to maximize the computation efficiency. (2) CPUs and GPUs cannot effectively handle the token shuffling process of our dynamic token pruning. We design a specific FPGA kernel to handle the token shuffling in the middle of model inferences. (3) CPUs and GPUs need complicated processes to address work-load imbalance, whereas on FPGA, we can design customized hardware modules for balancing work-load.
@@
In this paper, we propose an algorithm-hardware codesign for accelerating ViT inference. Different from existing ViT acceleration works, we utilize the combined power of static weight pruning and dynamic token pruning. We propose a simultaneous pruning algorithm to recover the model accuracy caused by two pruning approaches. Combining the two pruning approaches leads to more severe computational irregularity. Therefore, we develop a customized data path and memory organization on FPGA to execute the pruned model efficiently. While existing ViT accelerators on FPGA [35, 37] can handle the irregular patterns after pruning, they target either weight pruning or token pruning, but not both. Therefore, none of the existing FPGA ViT accelerators can support our integrated simultaneous pruning approach. We summarize our main contributions below:
@@
  • We propose an algorithm-hardware codesign for efficient ViT inference based on FPGA. The design combines parameter (static) and token (dynamic) pruning to reduce both the ViT model size and computational complexity.
@@
  • For the hardware design, we develop a novel hardware accelerator with multi-level parallelism and a load balancing strategy. This can efficiently deal with (1) load imbalance caused by the block pruning and (2) a changing number of tokens caused by token pruning. We also develop an efficient hardware mechanism for executing the on-the-fly token pruning algorithm.
  • We evaluate our codesign on DeiT models and deploy the proposed accelerator on a state-of-the-art FPGA board - Xilinx Alveo U250. The evaluation results show that the proposed algorithm can reduce computation complexity by up to 3.4×3.4$\times$3.4 × with $\approx 3\%$ accuracy drop and a model compression ratio of up to 1.6×1.6$\times$1.6 ×. Compared with state-of-the-art implementation on CPU, GPU, and FPGA, our codesign on FPGA achieves average latency reduction of 12.8×12.8$\times$12.8 ×, 3.2×3.2$\times$3.2 ×, 0.72.1×0.7-2.1$\times$0.7 - 2.1 × respectively.
@@ II-A Vision Transformer
ViT [1] has a stack of transformer encoders. Each encoder has a multi-head self-attention (MSA) and a multi-layer perceptron (MLP). The input image $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$ is first partitioned into $N$ patches $\mathbf{x}_{p}\in\mathbb{R}^{N\times P^{2}C}$. Each patch is flattened into a vector of length $P^{2}C$. Next, a learnable linear mapping method maps each patch to a token vector of length $D$. A special parameterized token $\mathbf{x}_{\text{CLS}}$ is appended as a token vector. Then, a positional embedding $\mathbf{E}_{\text{POS}}\in\mathbb{R}^{(N+1)\times D}$ is added to input token matrix to produce $\mathbf{Z_{0}}\in\mathbb{R}^{(N+1)\times D}$ which is the input to the transformer encoder stack. For simplicity, we denote the number of input tokens to the encoder stack as $N$ instead of $N+1$ for the rest of the paper.
@@ II-B Related Work
Weight pruning: Structured and hardware-friendly model parameter pruning, used in traditional CNNs [20, 21, 22], becomes popular for ViT. [17] introduces the notion of movement pruning, which prunes parameters by generating a pruning mask based on the learned scores. [23] proposes to prune parameters across all the encoders. Magnitude-based approaches, on the other hand, discard parameters with large magnitudes [24]. [18] partitions a parameter matrix into blocks and prunes the rows and columns in each block by using the $l_{2}$ norms. [25] prunes the entire attention heads within the MSA and neurons in each feed-forward linear layer (width pruning). [25] also removes entire encoders after a certain depth (depth pruning). [26] proposes a collaborative approach to optimizing ViT pruning that prunes heads and neurons; the latter neurons are pruned such that they reduce the length of each token. This method considers the collective pruning impact through an expensive approximation of the Hessian of the loss.
Token pruning: Token pruning approaches attempt to identify redundant tokens and drop them to reduce the computational footprint associated with the number of tokens. Both [19] and [39] have been notable for accelerating transformer models by leveraging the inherent sparsity in attention mechanisms. However, they do not use weight and token pruning simultaneously. [28] proposes a static approach to token dropping that ranks the importance of tokens by the attention score of the class token with respect to each token aggregated across heads. In theory, such static approaches do not need additional training since the token-dropping module is not parameterized. In contrast to static token pruning, dynamic token pruning as employed in [36, 35, 34] adds additional model parameters that facilitate the selection of relevant/attentive tokens. [36] and [35] utilize a token selector network inserted at various depths along the original transformer network; such token selector networks are neural networks that output a decision (binary) mask to inform the retention or removal of a token. [34], on the other hand, associate a learnable score to each token and prune tokens with scores lower than a threshold.
@@ III-A Problem Definition
Our objective is to accelerate ViT on FPGA through an algorithm-hardware codesign that 1) utilizes a novel combination of model weight pruning and token pruning to reduce computation complexity (algorithm design), and 2) an efficient accelerator that explicitly accounts for the distinct and irregular access patterns of the two pruning techniques for executing the combined pruned model (hardware design).
For the algorithm design, the input is a ViT model denoted as $\mathcal{M}(\cdot,h_{\text{structure}},\Theta)$ where $h_{\text{structure}}$ are the model hyper-parameters, including the number of encoders, number of heads, dimensions of linear layers, etc. $\Theta$ is the trainable parameters containing the weights and biases of the MSA and MLP. Our algorithm design prunes the input model $\mathcal{M}$ through (1) offline weight pruning that reduces the redundant parameters in $\Theta$, and (2) runtime token pruning that trims the number of tokens (in $h_{\text{structure}}$) in the intermediate layers according to the importance of the token. After pruning, we obtain the pruned model denoted as $\mathcal{M}^{{}^{\prime}}(\cdot,h_{\text{structure}}^{{}^{\prime}},\Theta^{{}^{\prime}})$, where $\Theta^{{}^{\prime}}$ denotes the model parameters after weight pruning and $h_{\text{structure}}^{{}^{\prime}}$ denotes the hyperparameters after token pruning. Our algorithm design aims to reduce the number of parameters, reduce the computation complexity, and maintain accuracy.
For the hardware design, the hardware accelerator executes the pruned model $\mathcal{M}^{{}^{\prime}}(\cdot,h_{\text{structure}}^{{}^{\prime}},\Theta^{{}^{\prime}})$. For executing the model, the input is an image sample $\bm{x}$, and the accelerator executes $\mathcal{M}^{{}^{\prime}}(\bm{x},h_{\text{structure}}^{{}^{\prime}},\Theta^{{}^{\prime}})$ to obtain the result. The latency is the duration from the time when the accelerator receives the input $\bm{x}$ to the time when the accelerator obtains the result.
@@ III-B Computational Complexity
The computational complexity for each operation within the MSA and MLP without pruning are summarized in table I. $B$ denotes the batch size.
@@ III-C Overview of Algorithm-hardware Codesign
Hardware accelerator design: At runtime, when the host process receives an input image, it sends the input image to the accelerator for inference. We employ a multi-level parallelism strategy for the proposed hardware architecture to efficiently handle the irregular computation patterns. We design a Token Dropping Hardware Module for efficient on-the-fly token dropping. See Section (Section V-C) for details.
@@ IV Pruning Algorithm
Existing works only utilize either weight pruning or token pruning. In contrast, we systematically combine the two pruning approaches with a novel training algorithm for recovering accuracy. To this end, we first introduce static weight pruning (Section IV-A) and dynamic token pruning (Section IV-B) separately, and then introduce our Simultaneous Pruning algorithm to prune the input model.
@@ IV-A1 Pruning of MSA
We use $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, $\mathbf{W}_{v}$ $\in\mathbb{R}^{D\times HD^{\prime}}$ to denote the concatenation of weight matrices of all the heads. For example, $\mathbf{W}_{p}\in\mathbb{R}^{D\times HD^{\prime}}\text{ }\mbox{where}\text{ }p=\{q,k,v\}.$ The projection operation (Equation 5) projects the concatenated SA outputs of embedding dimension $HD^{\prime}$ back to dimension $D$ via $\mathbf{W}_{\text{proj}}\in\mathbb{R}^{HD^{\prime}\times D}$. To prune a weight matrix $\mathbf{W}\in\mathbb{R}^{M_{1}\times M_{2}}$, we define a parameterized score matrix $\mathbf{S}\in\mathbb{R}^{m\times n}$ such that, $(m,n)=(\left\lceil\frac{M_{1}}{b}\right\rceil,\left\lceil\frac{M_{2}}{b}\right\rceil)$ where $(b,b)$ is the block size. $\mathbf{S}_{ij}$ denotes the importance score of a parameter block of size $(b,b)$ in the weight matrix $\mathbf{W}$ denoted by the slice $\mathbf{W}(ib:\alpha,jb:\beta)$ where $(\alpha,\beta)=(\min(ib+b,M_{1}),\min(jb+b,M_{2}))$. $\mathbf{S}$ is used to construct a mask $\mathbf{M}\in\mathbb{R}^{M_{1}\times M_{2}}$ via the top-$k$ selection:
@@
wher $\mathbf{M}_{ij}^{\text{block}}(.)$ is a block of size $(b,b)$ in $\mathbf{M}$ corresponding to $s_{ij}$. The masked weight is generated as, $\mathbf{W(\mathbf{M})}=\mathbf{W}\odot\mathbf{M}$ where $\odot$ is the element-wise Hadamard product. The generated masked weight $\mathbf{W(\mathbf{M})}$ is used for the forward pass during training. Note that top-$k$ is the target weight blocks of interest. To compute the gradient of $\mathbf{S}$ during the backward pass, a straight-through estimator (STE) [40, 41, 42] is used that neglects the gradients of $\mathbf{M}$ with respect to $\mathbf{S}$. Additionally, the pruning of $\mathbf{W}_{p}$ in row dimension and the pruning of $\mathbf{W}_{\text{proj}}$ in column dimension follows the same pattern (denoted as alternate pattern), as shown in Figure 2. For example, a head removed from $\mathbf{W}_{p}$ makes the corresponding head in $\mathbf{W}_{\text{proj}}$ redundant, and vice-versa.